OpenStudy (anonymous):

How do I eliminate the fractions in this equation? 7x/3+5y/4=4 5x/6-2y=21

OpenStudy (phi):

multiply by the denominator. for example, 3*4= 12 will get rid of the 3 and the 4 in the first equation.

OpenStudy (anonymous):

for the second one do i multiply the denominaotr by 6

OpenStudy (phi):

yes, because 6* (3x/6)= 3x and the denominator is eliminated.

OpenStudy (phi):

you should get 28x +15y= 48 5x -12y= 126

OpenStudy (phi):

to solve, multiply the top equation by 4 and the bottom by 5 then add the two equations

OpenStudy (anonymous):

why 4 and 5?

OpenStudy (anonymous):

i got x=6

OpenStudy (anonymous):

y=-8

OpenStudy (phi):

I chose to get rid of y (you could choose x, it does not matter much) to get rid of y, you want the same number (coefficient) in front of the y the top y term is 15y 15 is 3*5 bottom y term is -12y 12 is 3*4 least common multiple is 3*4*5= 60. we want both terms to have a 60 in front. so 4 for the top , 5 for the bottom

OpenStudy (phi):

yes, you can check by plugging those values into the original equations. But you did it correctly.

OpenStudy (anonymous):

thank you!